# Towards Pervasive Distributed Agentic Generative AI - A State of The Art

### v1

@@ 2.3.3. LLM-Evaluators
A growing number of researchers are also exploring the use of Large Language Models (LLMs) themselves as intermediaries for agent assessment. For instance, ALI-Agent (Wang et al., 2024e) leverages autonomous LLM-driven agents to automatically generate realistic test scenarios and iteratively refine them, enabling adaptive evaluations of alignment that effectively identify subtle, long-tail risks without relying on continuous human feedback. This framework employs a memory module for scenario generation, a tool-using module that integrates Web search and fine-tuned evaluators to reduce human labor, and an action module for test refinement. Building upon the concept of LLMs as judges, the Agent-as-a-Judge framework (Zhuge et al., 2024) extends this by employing agents to evaluate other agents. Recognizing the step-by-step operation of the agents, this framework aims to provide rich, intermediate feedback throughout the task-solving process, rather than relying solely on final outcomes. This approach has demonstrated success in code generation tasks, outperforming traditional LLM-as-a-Judge methods and achieving reliability comparable to human evaluations. ChatEval (Chan et al., 2023) employs a multi-agent debate approach to enhance LLM-based evaluation quality. By deploying a team of LLM agents with diverse role prompts, ChatEval facilitates autonomous discussion and evaluation of generated responses. This synergistic approach, leveraging the unique capabilities of multiple LLMs, exhibits superior accuracy and correlation with human assessments compared to single-agent evaluations. Similar multi-agent systems, such as IntellAgent (Levi and Kadar, 2025), use an LLM to perform roles like event generation, user agent, and dialog critique, implicitly integrating LLMs into the evaluation of agent behavior. SCALEEVAL (Chern et al., 2024) proposes an agent-debate-assisted meta-evaluation framework to address this, employing communicative LLM agents for iterative discussions that assist human annotators in identifying the most capable LLM evaluators, particularly in novel, user-defined scenarios. In cases of agent disagreement, minimal human oversight ensures a balance between efficiency and reliability. This approach enables scalable assessment of LLM evaluator trustworthiness across diverse tasks and criteria. While LLM-based evaluation offers advantages such as automation, scalability, and the ability to probe complex behaviors, it also presents challenges. These include reliance on inherent LLM capabilities and potential biases, which may cause incorrect evaluations, diverging from expert human judgment in domain-specific tasks (Szymanski et al., 2025). Furthermore, LLM evaluations are susceptible to prompt engineering and format variations, potentially leading to inconsistent assessments (Chern et al., 2024). The inherent lack of understanding in specialised domains can result in potential errors that human experts would make (Szymanski et al., 2025).
@@ 2.4. Field of Use
  • Software Engineering and Code Generation: Agents simulate developer workflows using tools like terminals and GitHub. Platforms like OpenHands (Wang et al., 2024b) enable secure execution of scripts and commands in sandboxed environments, supporting tool reuse and multi-agent collaboration. Reflexion (Shinn et al., 2023) enhances agent reasoning through verbal feedback and episodic memory. MAGIS (Tao et al., 2024) focuses on ongoing software maintenance using multiple LLM agents for bug fixes, feature updates, and optimizations.
  • Embodied Agents: These agents interact with physical or virtual environments using sensory feedback. LAC (Xu et al., 2024c) integrates LLMs into robotic control. Steve (Zhao et al., 2024a) and Voyager (Wang et al., 2023) operate in Minecraft environment, while WebPilot (Zhang et al., 2024e) and AutoWebGLM (Lai et al., 2024) automate web-based tasks.
  • Research and Development: Agents assist or automate the research process. Agent Laboratory (Schmidgall et al., 2025) supports literature review and experimentation. AI Scientist (Lu et al., 2024) enables full-cycle scientific discovery. ADAS (Hu et al., 2024) automates agent system design, allowing transfer across domains.
  • Information Management and Retrieval: Agents go beyond static information extraction by autonomously optimizing when and how to retrieve and process data. Systems like A-MEM (Xu et al., 2025) enable contextual memory and adaptive learning. RAG pipelines are enhanced by multi-agent RL architectures (Chen et al., 2025b), while Search-o1 (Li et al., 2025) empowers reasoning through agentic search and document reasoning.
  • Healthcare and Medicine: Agents automate complex clinical and administrative tasks. ClinicalAgent (Yue et al., 2024) leverages external biomedical knowledge for trial analysis. Other systems (Gebreab et al., 2024) handle administrative processes like record retrieval, patient registration, billing, and appointment scheduling.
@@
The demonstrated versatility of LLM-based agents across various domains motivates their exploration within everyday environments like smartphones and IoT devices. Their deployment in pervasive computing settings offers solutions for complex decision automation, enhanced user interaction, and improved accessibility. This potential is supported by recent advances, including compact and efficient LLMs (Team, 2025; Yang et al., 2025a), increased memory in embedded systems, power-efficient processors, specialised hardware such as NPUs, and high-speed communication technologies like 5G. The following sections explore the pervasive computing landscape, the evolving role of AI within it, and the integration of LLM-based agents in pervasive environments.
@@ 3. Pervasive Computing: an Overview
Pervasive computing, also referred to as ubiquitous computing, was conceived in early seminal work as technology that would ’weave themselves into the fabric of everyday life until they are indistinguishable from it’ (Weiser, 1991). This concept aims the creation of environments densely populated with computing and communication capabilities, yet seamlessly integrated with human users, until the technology effectively becomes ”transparent” (Satyanarayanan, 2001; Saha et al., 2003). So while distributed systems and mobile computing enabled ”anytime, anywhere” access to information, though not always with guaranteed connectivity, pervasive computing fundamentally shifts towards an ”all the time, everywhere” presence, ensuring seamless access to computing whenever and wherever it is needed (Saha et al., 2003). Pervasive computing manifests in various devices, formats, and locations, ranging from resource-constrained sensors to high-performance servers, including cloud datacenters, mobile edge computing servers, mobile devices, TVs, wearables, sensors and embedded systems. These interconnected devices leverage various wireless communication technologies to enhance their capabilities while minimizing resource consumption, including battery power, memory, and CPU time leading to the proliferation of Internet of Things (IoT) devices. These IoT devices, endowed with sensing, computing, networking, and communication functionalities, are capable of collecting, analysing, and transmitting a diverse spectrum of data, including images, videos, audio, texts, wireless signals, and physiological signals from individuals and the physical environment. Furthermore sensors, integral to many IoT devices, play a crucial role by monitoring phenomena both within and beyond human perception, converting physical occurrences into numerical data. The size reduction, increased efficiency, enhanced sensitivity, and improved connectivity of these sensors have enabled their embedding in diverse environments and objects, providing real-time data and insights previously inaccessible. Therefore this increase in connectivity, data generation and accessibility of sensory data, contributes to the generation of zettabytes of real-time data streams (Baccour et al., 2022) that can be used to create much more customizable and accurate systems but also can be challenging to manage. These pervasive systems find application across diverse sectors, ranging from Smart Homes and Smart Cities to Smart Factory and Healthcare. Cisco’s projections [^1] Cisco 5G IoT White Paper indicate over 2.6 billion cellular-connected IoT devices by 2026, showing the maturation of their capabilities and decreasing costs. Furthermore, the economic impact of IoT is projected to be substantial, with estimates suggesting a global value creation ranging from \$5.5 trillion to \$12.6 trillion by 2030 (Chui et al., 2021). All these factors underscore the substantial and growing impact of pervasive computing environments in improving the quality of life.
@@ 3.1. Computational and Network Advancements
The increasing diffusion of pervasive computing is significantly driven by the advancements in both computational power and network technologies. Embedding computational power within everyday objects is fundamental to pervasive computing. This necessitates a range of microprocessor solutions tailored to specific device requirements. These range from low-power Microcontroller Units (MCUs) for basic sensing and actuation in resource-constrained devices, to more powerful Central Processing Units (CPUs) for complex processing tasks, and highly integrated Systems on a Chip (SoCs) for feature-rich applications. SoCs often incorporate specialised processing units such as Graphics Processing Units (GPUs) and Digital Signal Processors (DSPs) to enhance multimedia and signal processing capabilities within compact form factors. Notably, the emergence of architectures like PlasticARM CPUs (Biggs et al., 2021) presents a promising avenue for pervasive devices, including smartphones and edge computing nodes, offering an optimised balance of computational performance and energy efficiency crucial for their operation. This continuous evolution in microprocessor technology allows for increasingly sophisticated computational intelligence to be integrated directly into the fabric of everyday life. Also the increasing integration of Artificial Intelligence into edge devices within pervasive computing environments demands dedicated processing capabilities for computationally intensive AI tasks. Recognizing the limitations of general-purpose CPUs and GPUs for these specialised workloads, the development of dedicated AI accelerators has been a critical enabler. These specialised silicon solutions, often implemented as discrete co-processors or integrated within SoCs as Neural Processing Units (NPUs) (Tan and Cao, 2021), offer significantly enhanced performance and energy efficiency for neural network computations. The strategic inclusion of NPUs in SoCs designed for IoT, edge computing, and mobile platforms facilitates the rapid and efficient execution of complex AI algorithms directly on the device, minimizing latency and power consumption associated with cloud-based processing. The seamless operation of interconnected devices, a defining characteristic of pervasive computing, relies critically on robust and efficient networking infrastructure. Given the heterogeneity of devices and their diverse communication requirements, a suite of networking technologies is essential. While high-bandwidth applications benefit from advancements in Wi-Fi standards like Wi-Fi 7 with its focus on minimal latency, low-power devices leverage technologies such as ZigBee for personal area networks. Wide-area connectivity for distributed IoT deployments is facilitated by Low Power Wide Area Networks (LPWANs) like LoRaWAN (Bimpas et al., 2024). Furthermore, high-speed cellular technologies like 5G and 6G provide ubiquitous connectivity for mobile and edge devices. This diverse and evolving network infrastructure provides the essential connectivity for the complex communication and coordination within pervasive computing environments.
3.2. Architectural Infrastructures
Figure 4. The Cloud, Fog and Edge layers
@@ 3.2. Architectural Infrastructures
Pervasive computing leverages various architectural infrastructures to deliver its capabilities, primarily involving the cloud, fog, and edge computing paradigms (Figure 4). The cloud layer offers high computational power and storage for complex tasks and centralised management. The fog layer, situated closer to the edge, facilitates distributed, latency-aware applications by providing local computing and network connectivity. Finally, the edge layer includes the end devices themselves, enabling processing and decision-making at the data source for real-time interactions. These architectures work in a complementary way to enable the collection, processing, and use of data generated by a multitude of interconnected devices.
3.2.1. Cloud Computing
@@
As defined by the National Institute of Standards and Technology (NIST), cloud computing is ’a model for enabling ubiquitous, convenient, on-demand network access to a shared pool of configurable computing resources (e.g., networks, servers, storage, applications, and services) that can be rapidly provisioned and released with minimal management effort or service provider interaction’ (Mell et al., 2011). This paradigm includes centralised or distributed computing technologies operating over the Internet, primarily functioning as a scalable storage and processing infrastructure. Parallel and distributed computing models can be independently or jointly integrated and deployed within data centers, either physically or as virtualised resources. Within the context of pervasive computing, the cloud serves as a critical centralised repository and high-capacity processing hub for the extensive datasets generated by edge devices. Its inherent computational resources and scalability are paramount for managing high data flood (Yalli et al., 2024). Cloud service models are commonly categorised into private, community, public, and hybrid deployments (Mell et al., 2011). Private clouds are dedicated to a single organization, regardless of who manages it or its location. Community clouds serve a specific group with shared needs, potentially managed by members or a third party, on or off-site. Public clouds are openly accessible to anyone and are owned and operated by providers on their premises. Hybrid clouds combine two or more distinct cloud types, linked by technology allowing data and application movement between them. Thanks to cloud computing, the connectivity of IoT devices to the cloud, and their integration with other related sensors, unlocks significant potentials:
@@
3.2.2. Fog Computing
The key goal of the fog layer, is to reduce the gap between the cloud layer, which has high resources, and the edge layer, which has limited resources. As defined by the National Institute of Standards and Technology (NIST), fog computing is ’a layered model for enabling ubiquitous access to a shared continuum of scalable computing resources’ where ’facilitates the deployment of distributed, latency-aware applications and services, and consists of fog nodes (physical or virtual), residing between smart end-devices and centralised (cloud) services’ (Iorga et al., 2017). Fog nodes can be seen as physical components such as Raspberry Pi, Nividia Jetson platforms, gateways, switches, routers, nano-servers, or virtual components such as virtualised switches, virtual machines, or cloudlets, bringing network, storage, and computing capabilities closer to the end users, and offering several unique features including reduced latency, geographical distribution, enhanced data security, and real-time processing (Sabireen and Neelanarayanan, 2021; Yalli et al., 2024). Fog computing can handle large volumes of data produced by various edge device types instead of sending it to a central cloud infrastructure, thereby addressing bandwidth and energy consumption problems. It possesses the ability to handle large volumes of data (potentially better than the cloud in terms of energy consumption), process data quickly, and produce high-quality results. Edge devices are typically located in close proximity or within a short distance of the fog layer, ensuring faster communication between these two tiers. However, fog computing also presents several challenges, including:

### v2

@@ 2.3.3. LLM-Evaluators
A growing number of researchers are also exploring the use of Large Language Models (LLMs) themselves as intermediaries for agent assessment. For instance, ALI-Agent (wang2024ali) leverages autonomous LLM-driven agents to automatically generate realistic test scenarios and iteratively refine them, enabling adaptive evaluations of alignment that effectively identify subtle, long-tail risks without relying on continuous human feedback. This framework employs a memory module for scenario generation, a tool-using module that integrates Web search and fine-tuned evaluators to reduce human labor, and an action module for test refinement. Building upon the concept of LLMs as judges, the Agent-as-a-Judge framework (zhuge2024agent) extends this by employing agents to evaluate other agents. Recognizing the step-by-step operation of the agents, this framework aims to provide rich, intermediate feedback throughout the task-solving process, rather than relying solely on final outcomes. This approach has demonstrated success in code generation tasks, outperforming traditional LLM-as-a-Judge methods and achieving reliability comparable to human evaluations. ChatEval (chan2023chateval) employs a multi-agent debate approach to enhance LLM-based evaluation quality. By deploying a team of LLM agents with diverse role prompts, ChatEval facilitates autonomous discussion and evaluation of generated responses. This synergistic approach, leveraging the unique capabilities of multiple LLMs, exhibits superior accuracy and correlation with human assessments compared to single-agent evaluations. Similar multi-agent systems, such as IntellAgent (levi2025intellagent), use an LLM to perform roles like event generation, user agent, and dialog critique, implicitly integrating LLMs into the evaluation of agent behavior. SCALEEVAL (chern2024can) proposes an agent-debate-assisted meta-evaluation framework to address this, employing communicative LLM agents for iterative discussions that assist human annotators in identifying the most capable LLM evaluators, particularly in novel, user-defined scenarios. In cases of agent disagreement, minimal human oversight ensures a balance between efficiency and reliability. This approach enables scalable assessment of LLM evaluator trustworthiness across diverse tasks and criteria. While LLM-based evaluation offers advantages such as automation, scalability, and the ability to probe complex behaviors, it also presents challenges. These include reliance on inherent LLM capabilities and potential biases, which may cause incorrect evaluations, diverging from expert human judgment in domain-specific tasks (szymanski2025limitations). Furthermore, LLM evaluations are susceptible to prompt engineering and format variations, potentially leading to inconsistent assessments (chern2024can). The inherent lack of understanding in specialised domains can result in potential errors that human experts would make (szymanski2025limitations).
@@ 2.4. Field of Use
  • Software Engineering and Code Generation: Agents in this domain have moved beyond simple code completion to handling full development lifecycles. Current research focuses on automating distinct engineering tasks. First, agents perform secure environment interaction and workflow simulation; for instance, platforms like OpenHands (wang2024openhands) execute scripts and commands in sandboxed environments to mimic human developer workflows. Second, agents tackle iterative debugging and self-correction. For instance Reflexion (shinn2023reflexion) uses verbal feedback and episodic memory to reason through and fix code errors or MAGIS (tao2024magis), which employs a multi-agent framework to resolve GitHub issues, implement feature updates, and perform code optimizations.
  • Embodied Agents: In physical and simulated environments, agents are tasked with grounding high-level language into low-level control policies. Recent approaches leverage foundation models to bridge this gap across distinct domains. In robotic control, Magma (yang2025magma) introduces spatial-temporal reasoning via Set-of-Mark and Trace-of-Mark to bridge verbal and spatial intelligence, while LAC (xu2024lac) employs a multi-agent controller framework that merges LLM reasoning with feedback control loops for real-time responsiveness. For open-ended exploration, agents are used in Minecraft sandboxes: Voyager (wang2023voyager) demonstrates lifelong learning through an automatic curriculum and an executable code skill library, whereas Steve (zhao2024see) enhances planning by integrating visual perception components with language instruction. Finally agents can be used in web task execution. For instance WebPilot (zhang2024webpilotversatileautonomousmultiagent) addresses dynamic environments through a dual-strategy Monte Carlo Tree Search (MCTS), while AutoWebGLM (lai2024autowebglm) uses simplified HTML representations and reinforcement learning to navigate complex real-world webpages effectively.
  • Research and Development: Agents are increasingly tasked with automating distinct phases of the scientific method. For lifecycle management, frameworks like Agent Laboratory (schmidgall2025agent) act as interactive co-pilots, coordinating specialized sub-agents to handle literature reviews, experimentation, and report generation. In the field of autonomous discovery, the AI Scientist Agent (lu2024ai) pushes boundaries by generating novel ideas, executing code-based experiments, and authoring full manuscripts without human intervention. Also ADAS (hu2024automated) iteratively programs and optimizes new agentic architectures in code space, discovering novel building blocks with cross-domain transferability.
  • Information Management and Retrieval: Agents go beyond static information extraction by autonomously optimizing when and how to retrieve and process data. They are increasingly used for collaborative pipeline optimization, where individual RAG components (e.g., query rewriters, selectors) operate as cooperative agents trained via Multi-Agent Reinforcement Learning to maximize global answer quality (chen2025improving). Additionally, agents facilitate dynamic knowledge reasoning, exemplified by frameworks that autonomously trigger searches during chain-of-thought processes and employ dedicated modules to refine retrieved documents into logic-compatible insights (li2025search).
  • Healthcare and Medicine: In critical safety domains, agents automate complex clinical and administrative tasks. For clinical trial optimization, agents like ClinicalAgent (yue2024clinicalagent) utilize external biomedical knowledge graphs to predict trial outcomes by autonomously assessing drug efficacy, safety risks, and enrollment feasibility. Agents are also used for hospital administration, for istance multi-agent frameworks (gebreab2024llm) streamline operations by autonomously navigating Electronic Medical Record (EMR) platforms to execute tasks such as patient registration, billing, and appointment scheduling.
@@
The versatility of LLM-based agents demonstrated across various domains motivates their exploration within everyday environments like smartphones and IoT devices. Their deployment in pervasive computing settings offers solutions for complex decision automation, enhanced user interaction, and improved accessibility. This potential is supported by recent advances, including compact and efficient LLMs (qwq32b; yang2025magma), increased memory in embedded systems, power-efficient processors, specialised hardware such as NPUs, and high-speed communication technologies like 5G. The following sections explore the pervasive computing landscape, the evolving role of AI within it, and the integration of LLM-based agents in pervasive environments.
@@ 3. Pervasive Computing: an Overview
Pervasive computing, also referred to as ubiquitous computing, was conceived in early seminal work as technology that would ’weave themselves into the fabric of everyday life until they are indistinguishable from it’ (weiser1991computer). This concept focuses on the creation of environments densely populated with computing and communication capabilities, yet seamlessly integrated with human users, until the technology effectively becomes ”transparent” (satyanarayanan2001pervasive; saha2003pervasive). So while distributed systems and mobile computing enabled ”anytime, anywhere” access to information, though not always with guaranteed connectivity, pervasive computing fundamentally shifts towards an ”all the time, everywhere” presence, ensuring seamless access to computing whenever and wherever it is needed (saha2003pervasive). Pervasive computing manifests in various devices, formats, and locations, ranging from resource-constrained sensors to high-performance servers, including cloud datacenters, mobile edge computing servers, mobile devices, TVs, wearables, sensors and embedded systems. These interconnected devices leverage various wireless communication technologies to enhance their capabilities while minimizing resource consumption, including battery power, memory, and CPU time leading to the proliferation of Internet of Things (IoT) devices. These IoT devices, endowed with sensing, computing, networking, and communication functionalities, are capable of collecting, analysing, and transmitting a diverse spectrum of data, including images, videos, audio, texts, wireless signals, and physiological signals from individuals and the physical environment. Furthermore sensors, integral to many IoT devices, play a crucial role by monitoring phenomena both within and beyond human perception, converting physical occurrences into numerical data. The size reduction, increased efficiency, enhanced sensitivity, and improved connectivity of these sensors have enabled their embedding in diverse environments and objects, providing real-time data and insights previously inaccessible. Therefore this increase in connectivity, data generation and accessibility of sensory data, contributes to the generation of zettabytes of real-time data streams (baccour2022pervasive) that can be used to create much more customizable and accurate systems but also can be challenging to manage. These pervasive systems find application across diverse sectors, ranging from Smart Homes and Smart Cities to Smart Factory and Healthcare. Cisco’s projections [^1] Cisco 5G IoT White Paper indicate over 2.6 billion cellular-connected IoT devices by 2026, showing the maturation of their capabilities and decreasing costs. Furthermore, the economic impact of IoT is projected to be substantial, with estimates suggesting a global value creation ranging from \$5.5 trillion to \$12.6 trillion by 2030 (chui2021internet). All these factors underscore the substantial and growing impact of pervasive computing environments in improving the quality of life.
@@ 3.1. Computational and Network Advancements
The increasing diffusion of pervasive computing is significantly driven by the advancements in both computational power and network technologies. Embedding computational power within everyday objects is fundamental to pervasive computing. This necessitates a range of microprocessor solutions tailored to specific device requirements. These range from low-power Microcontroller Units (MCUs) for basic sensing and actuation in resource-constrained devices, to more powerful Central Processing Units (CPUs) for complex processing tasks, and highly integrated Systems on a Chip (SoCs) for feature-rich applications. SoCs often incorporate specialised processing units such as Graphics Processing Units (GPUs) and Digital Signal Processors (DSPs) to enhance multimedia and signal processing capabilities within compact form factors. Notably, the emergence of architectures like PlasticARM CPUs (biggs2021natively) presents a promising avenue for pervasive devices, including smartphones and edge computing nodes, offering an optimised balance of computational performance and energy efficiency crucial for their operation. This continuous evolution in microprocessor technology allows for increasingly sophisticated computational intelligence to be integrated directly into the fabric of everyday life. Also the increasing integration of Artificial Intelligence into edge devices within pervasive computing environments demands dedicated processing capabilities for computationally intensive AI tasks. Recognizing the limitations of general-purpose CPUs and GPUs for these specialised workloads, the development of dedicated AI accelerators has been a critical enabler. These specialised silicon solutions, often implemented as discrete co-processors or integrated within SoCs as Neural Processing Units (NPUs) (tan2021efficient), offer significantly enhanced performance and energy efficiency for neural network computations. The strategic inclusion of NPUs in SoCs designed for IoT, edge computing, and mobile platforms facilitates the rapid and efficient execution of complex AI algorithms directly on the device, minimizing latency and power consumption associated with cloud-based processing. The seamless operation of interconnected devices, a defining characteristic of pervasive computing, relies critically on robust and efficient networking infrastructure. Given the heterogeneity of devices and their diverse communication requirements, a suite of networking technologies is essential. While high-bandwidth applications benefit from advancements in Wi-Fi standards like Wi-Fi 7 with its focus on minimal latency, low-power devices leverage technologies such as ZigBee for personal area networks. Wide-area connectivity for distributed IoT deployments is facilitated by Low Power Wide Area Networks (LPWANs) like LoRaWAN (bimpas2024leveraging). Furthermore, high-speed cellular technologies like 5G and 6G provide ubiquitous connectivity for mobile and edge devices. This diverse and evolving network infrastructure provides the essential connectivity for the complex communication and coordination within pervasive computing environments.
@@
Pervasive computing leverages various architectural infrastructures to deliver its capabilities, primarily involving the edge-to-cloud continuum paradigm (Figure LABEL:fig:pervasive_layer). The cloud layer offers high computational power and storage for complex tasks and centralised management. The edge layer, situated closer to the user and the data, facilitates distributed, latency-aware applications by providing local computing, network connectivity and decision-making at the data source for real-time interactions. These architectures work in a complementary way to enable the collection, processing, and use of data generated by a multitude of interconnected devices.
3.1.1. Cloud Layer
@@
As defined by the National Institute of Standards and Technology (NIST), cloud computing is ’a model for enabling ubiquitous, convenient, on-demand network access to a shared pool of configurable computing resources (e.g., networks, servers, storage, applications, and services) that can be rapidly provisioned and released with minimal management effort or service provider interaction’ (mell2011nist). This paradigm includes centralised or distributed computing technologies operating over the Internet, primarily functioning as a scalable storage and processing infrastructure. Parallel and distributed computing models can be independently or jointly integrated and deployed within data centres, either physically or as virtualised resources. Within the context of pervasive computing, the cloud serves as a high-capacity processing hub for the extensive datasets generated by edge devices. Its inherent computational resources and scalability are paramount for managing high data flood (yalli2024internet). Cloud service models are commonly categorised into private, community, public, and hybrid deployments (mell2011nist). Private clouds are dedicated to a single organization, regardless of who manages it or its location. Community clouds serve a specific group with shared needs, potentially managed by members or a third party, on or off-site. Public clouds are openly accessible to anyone and are owned and operated by providers on their premises. Hybrid clouds combine two or more distinct cloud types, linked by technology allowing data and application movement between them. Thanks to cloud computing, the connectivity of IoT devices to the cloud, and their integration with other related sensors, unlocks significant potentials:
@@
3.1.2. Edge Layer
This is the pivotal layer of the pervasive architecture, performing computations closer to the data source to enhance responsiveness and efficiency. As defined by the National Institute of Standards and Technology (NIST), edge computing ’is the network layer including the end-devices and their users, to provide, for example, local computing capability on a sensor, metering or some other devices that are network-accessible’ (iorga2017nist). The edge layer has a wide variety of heterogeneous devices, spanning from the mere sensors and actuators, to resource-constrained devices (e.g., microcontrollers) capable of basic logic, up to powerful user-end units (e.g., smartphones, PCs) that can provide high computational capabilities (sun2016edgeiot; baccour2022pervasive). By decentralizing intelligence and processing power, this paradigm leverages this entire spectrum situated in close proximity to the data-generating entity (yalli2024internet).
To bridge the gap between these distributed edge devices and the centralized cloud, the architecture incorporates an intermediate infrastructure layer, historically referred to as Fog Computing. Functioning as a high-performance tier within the edge continuum, these nodes reside physically between smart end-devices and centralized services to facilitate connectivity and heavy processing. As defined by NIST, this infrastructure ’facilitates the deployment of distributed, latency-aware applications and services, and consists of fog nodes (physical or virtual), residing between smart end-devices and centralised services’ (iorga2017nist). These nodes include physical components such as Raspberry Pi, NVIDIA Jetson platforms, gateways, routers, or nano-servers, which are capable of handling larger volumes of data than simple sensors while avoiding the latency of the cloud (sabireen2021review; yalli2024internet). This intermediate tier usefull for offloading computations from resource-constrained edges, addressing bandwidth and energy consumption problems by processing data locally to produce high-quality results before transmission.
A specific implementation of the Edge Layer is Mobile Edge Computing (MEC), which deploys services and computational capabilities at the edge of cellular networks, close to subscribers. This strategic placement enables a service environment characterised by ultra-low latency, high bandwidth, and direct access to real-time network information (li2016mobile). For instance, in (sun2016edgeiot), researchers handle data streams at the mobile edge to overcome the scalability limits of traditional IoT architectures, reducing traffic load in the core network.
However, implementing and managing this heterogeneous edge environment presents several challenges: